When a gas is bubbled through water at 298K a very dilute solution of the gas is obtained. Henry's law of constant for the gas at 298K is 150K bar. If the gas excerts a partial pressure of 2 bar the no. Of millimoles of the gas dissolved in 1L of water is


1
Expert's answer
2021-10-04T05:09:14-0400

Henry’s law:

"P =K_Hx_{gas} \\

2 = 150 \times 10^3 \times x \\

x = \frac{2}{150 \times 10^3} = 0.0133 \times 10^{-3} \;moles \\

x = 0.0133 \;millimoles"